OpenStudy (amistre64):

what is the equation of the tangent plane to f(x,y) = x^2-y^2 +9 at the point (4,5,0) ?

OpenStudy (anonymous):

Razzleberry squared?

OpenStudy (amistre64):

close :)

OpenStudy (anonymous):

No? How about... 3 Penguins and a goose?

OpenStudy (amistre64):

it begins with a gradient :)

OpenStudy (anonymous):

A+.

OpenStudy (amistre64):

x^2-y^2-z+9=0 g=<2x,2y,-1> p=( 4, 5, 0) ----------- n= <8,10,-1>

OpenStudy (anonymous):

Oh so close!

OpenStudy (amistre64):

8(x-4)+10(y-5)-z=0 right?

OpenStudy (anonymous):

Hmm yes

OpenStudy (amistre64):

now you aint agreeing with me just cause im right are ya?

OpenStudy (anonymous):

Why else would would I?

OpenStudy (amistre64):

i like this purpley pic; its the moon by a streetlamp about dusk time

OpenStudy (amistre64):

just before the mosquitoes come out ;)

OpenStudy (amistre64):

what you know about double integrals?

OpenStudy (anonymous):

You have to take the partial derivative with respect to x and then with respect to y....so...fx=2x and fy=2y...solving this with the given points you can figure it out. So, fx(4)=8 and fy(5)=10...so...8(x-4)+10(y-5)-z=0......Ya!...I believe that's correct sir!

OpenStudy (anonymous):

ur playing again amistre..tut..tut

OpenStudy (amistre64):

...hangs head in shame.....but mr owletan, can i have another bwol of swill please?